CHM 2210 Homework 14 I have GOT to get me one of these! C.1. What is the major organic product from the reaction of 2-methy stuck
Answer #1

1) The reaction follows markovnikov's addition , that is negative part of the reagent goes to the carbon containing least number of hydrogen atom.

CH3 СН2 CH3 – C=CH-CH3 + H By – CH3-C- CH2 - CH3 2-methyl-2- Butene 2-Bromo-2-methylbutane

so the major product is 2-Bromo-2-methyl butane.

2) This reaction follows anti Markovnikov's addition, here the negative part of the reagent goes to the carbon containing more number of hydrogen atoms.

CH3 CH3-C = CH-CH3 H₂ Oz - + H By 2 2-methyl-2- Butene CH3-C- CH - CH₂ H Br 2-Bromo - 3-methyl butane

3) Mechanism

Mechanism of Reaction (1) [Mar kounikovs addition] Hor V HBT → 2-methyl 2-Butene 2- bromo-2-methyl butane Mechanism of React

4) The first reaction acts as an ionic reagent and it is the Br- reacts with the double bond of the alkene.

The second reaction is a radical reaction. Hydrogen peroxide acts as the radical initiator.

The first reaction follows markovnikov's addition , that is negative part of the reagent goes to the carbon containing least number of hydrogen atom.

The second reaction follows anti Markovnikov's addition, here the negative part of the reagent goes to the carbon containing more number of hydrogen atoms.

5) 1 +420 H2SO4, r ton 3,3-dimethyl-1- butene 2,3-dimethyl-2-butanol